### v1

# Interaction Screening and Pseudolikelihood Approaches for Tensor Learning in Ising Models

Abstract.
In this paper, we study two well known methods of Ising structure learning, namely the pseudolikelihood approach and the interaction screening approach, in the context of tensor recovery in $k$-spin Ising models. We show that both these approaches, with proper regularization, retrieve the underlying hypernetwork structure using a sample size logarithmic in the number of network nodes, and exponential in the maximum interaction strength and maximum node-degree. We also track down the exact dependence of the rate of tensor recovery on the interaction order $k$, that is allowed to grow with the number of samples and nodes, for both the approaches. Finally, we provide a comparative discussion of the performance of the two approaches based on simulation studies, which also demonstrate the exponential dependence of the tensor recovery rate on the maximum coupling strength.
Key words and phrases:
structure learning, Ising model, tensor, interaction screening, pseudolikelihood
1. Introduction
The $k$-spin Ising model [4, 35, 20, 9, 32] is a discrete exponential family for modeling dependent binary (±$\pm 1$-valued) data exhibiting multi-body interactions, which can be thought of as taking place along the hyperedges of a $k$-uniform hypergraph. It is a generalization of the classical $2$-spin Ising model, that was originally introduced by physicists for modelling ferromagnetism [26], and has since then found extensive applications in diverse areas such as social sciences, image processing, computational biology, neural networks, spatial statistics and election forecasting [3, 18, 23, 22, 25, 31, 27, 28]. However, pairwise interactions are often not adequate to capture all the complex dependencies that arise in real-world network structures, such as peer-group effects and multi-atomic interactions on crystal surfaces, which necessitates the use of higher order tensor Ising models for modeling such complex relational frameworks. The multi-body interactions in tensor Ising models are captured by an interaction tensor, which in many practical cases, forms the adjacency tensor of an underlying hypergraph, whose nodes are thought of as being sites for binary-valued variables or spins, that interact with each other along the hyperedges of this hypergraph. A fundamental problem in the area of Ising model learning is to recover this underlying interaction structure given access to multiple samples from this model, on which a significant amount of literature has developed for the classical $2$-spin Ising models, during the past two decades (see, for example, [2, 10, 16, 24, 36, 37, 30]). Variants of the complete structure learning problem have also experienced growing interest over the past few years, some related notable works being graph property testing by Neykov et al. [14, 34], identity and independence testing by Daskalakis et al. [17], and Ising model learning by Devroye et al. [19].
Ravikumar et al. [36] implemented an $\ell_{1}$-penalized nodewise logistic regression approach to recover the signed neighborhood of each node of the underlying Ising graph. However, as pointed out in [5], the incoherence condition assumed in [36] is $\mathrm{NP}$-hard to check for general non-ferromagnetic Ising models, and is computationally expensive in the ferromagnetic case, too. Bresler [10] proposed a greedy algorithm for exact recovery of an Ising interaction matrix using a sample of size logarithmic in the network size, that has runtime complexity much lower than the $\ell_{1}$-regularized nodewise logistic regression approach in [36]. However, practically speaking, the applicability of this method does not extend much beyond bounded-degree networks, since the dependence of the sample size on the maximum node degree is double exponential in nature. For related works on structure learning in antiferromagnetic and tree-structured Ising models, see [11, 12]. Vuffray et al. [38] proposed a quasi-quartic time algorithm based on interaction screeing, that also works on sample sizes logarithmic in the network size, but exponential in the order of the maximum node degree and the maximum interaction strength. This scaling, unlike the algorithm in [10], matches the information theoretic lower bound of [37]. As an added advantage, unlike the approach in [10], the interaction screening approach in [38] does not require prior knowledge of model parameters, such as the maximum interaction strength and largest node degree, which may not always be obtainable.
In spite of the presence of a significant amount of literature on matrix structure learning in classical $2$-spin Ising models, to the best of our knowledge, analogous literature for tensor recovery in $k$-spin Ising models is extremely scarce, if at all existent. Only recently, Liu at al. [29] extended the technique and results in [36] to tensor recovery in $k$-spin Ising models. In this paper, we study the interaction screening [38] and the regularized pseudolikelihood methods for tensor recovery in $k$-spin Ising models. As was observed in [5], the regularized pseudolikelihood algorithm performs very poorly for high values of the maximum coupling strength, despite working with a sample size much larger than the logarithm of the number of nodes, and irrespective of the choice of the LASSO tuning parameter. Further, the threshold for the maximum coupling strength, above which this poor performance emerges, is also very near to the critical point for the Ising model on a randomly diluted grid (see [5]). The dependence of the minimum sample size on the maximum coupling strength was not made explicit in [36] or [29], and was incorporated within the proportionality constant. Inspired by [38], in this paper, we work out this dependency for tensor Ising models, and show that a sample size exponential in the maximum coupling strength guarantees consistent tensor recovery. This is a plausible explanation behind the worse behavior of the regularized estimators at low temperatures (high coupling strengths). We also track down the dependence of the minimum sample size and the rate of convergence on the order $k$ of interaction. In the following subsection, we describe the $k$-spin Ising model and the tensor recovery problem in more details.
1.1. The Tensor Recovery Problem
The $k$-spin or $k$-tensor Ising model is a probability distribution on the set $\{-1,1\}^{p}$, defined as:
${\mathbb{P}}_{\bm{J}}(\bm{x}):=\frac{1}{Z(\bm{J})}e^{H(\bm{x})}\quad(\bm{x}\in\{-1,1\}^{p})$ (1)
where $\bm{J}:=((J_{r_{1},\ldots,r_{k}}))_{(r_{1},\ldots,r_{k})\in[p]^{k}}$ denotes a $k$-fold tensor with $[p]:=\{1,\ldots,p\}$,
$H(\bm{x}):=\sum_{(r_{1},\ldots,r_{k})\in[p]^{k}}J_{r_{1},\ldots,r_{k}}x_{r_{1}}\ldots x_{r_{k}},$
and $Z(\bm{J})$ is a normalizing constant required to ensure that ${\mathbb{P}}_{\bm{J}}$ is a valid probability distribution. Hereafter, we will assume that the tensor $\bm{J}$ satisfies the following properties:
  1. (1)
    $\bm{J}$ is symmetric, i.e., $J_{r_{1},...,r_{k}}=J_{r_{\sigma(1)},...,r_{\sigma(k)}}$ for every $(r_{1},\ldots,r_{k})\in[p]^{k}$ and every permutation $\sigma$ of $\{1,...,k\}$,
  2. (2)
    $\bm{J}$ has zeros on the diagonals, i.e., $\bm{J}_{r_{1}...r_{k}}=0$, if $r_{s}=r_{t}$ for some $1\leqslant s<t\leqslant k$.
Also, throughout the paper, we will denote the maximum entry of $\bm{J}$ by $\beta$, i.e.
$\beta:=\max_{(r_{1},\ldots,r_{k})\in[p]^{k}}J_{r_{1},\ldots,r_{k}}~{}.$
The aim of this paper is to recover the unknown tensor $\bm{J}$ given access to samples $\bm{x}^{(1)},\bm{x}^{(2)},\ldots,\bm{x}^{(n)}$ from this model. Based on the works [30, 38] in the context of Ising matrix recovery, in this paper, we analyze the performance of two approaches, namely the regularized interaction screening estimator and the regularized pseudolikelihood estimator on tensor recovery in higher-order Ising models.
The regularized interaction screening estimator (RISE) of the neighborhood of a particular node $r\in[p]$ is defined as:
$\hat{\bm{J}}_{r,I}:=\arg\min_{\bm{J}_{r}\in\mathbb{R}^{{\binom{p-1}{k-1}}}}\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})+\lambda\|\bm{J}_{r}\|_{1}$ (2)
where $\bm{J}_{r}:=((J_{r,r_{1},\ldots,r_{k-1}}))_{(r_{1},\ldots,r_{k-1})\in T_{r}}$ and
$\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n}):=\frac{1}{n}\sum_{i=1}^{n}\exp\left(-kx^{(i)}_{r}m_{r}(\bm{x}^{(i)})\right),$
with $m_{r}(\bm{x}):=\sum_{(r_{1},\ldots,r_{k-1})\in[p]^{k-1}}J_{r,r_{1},\ldots,r_{k-1}}x_{r_{1}}\ldots x_{r_{k-1}}$ and
$T_{r}:=\{(r_{1},\ldots,r_{k-1})\in([p]\setminus\{r\})^{k-1}:1\leq r_{1}<\ldots<r_{k-1}\leq p\}.$
Note that the objective function in (3) is convex, and hence, (3) is a convex optimization problem.
On the other hand, the regularized pseudolikelihood estimator (RPLE) of the neighborhood of the node $r\in[p]$ is defined as:
$\hat{\bm{J}}_{r,P}:=\arg\min_{\bm{J}_{r}\in\mathbb{R}^{{\binom{p-1}{k-1}}}}\ell(\bm{J}_{r};\mathfrak{X}^{n})+\lambda\|\bm{J}_{r}\|_{1}$ (3)
where
$\ell(\bm{J}_{r};\mathfrak{X}^{n}):=-\frac{1}{n}\sum_{i=1}^{n}\log{\mathbb{P}}_{\bm{J}}(x_{r}^{(i)}|\bm{x}_{\setminus r}^{(i)})$
and $\bm{x}_{\setminus r}^{(i)}:=(x_{t}^{(i)})_{t\neq r}$. A straightforward computation shows that:
${\mathbb{P}}_{\bm{J}}(x_{r}|\bm{x}_{\setminus r})=\frac{\exp(kx_{r}m_{r}(\bm{x}))}{2\cosh(kx_{r}m_{r}(\bm{x}))}$
and hence, one has:
$\ell(\bm{J}_{r};\mathfrak{X}^{n})=-\frac{1}{n}\sum_{i=1}^{n}\left\{kx_{r}^{(i)}m_{r}(\bm{x}^{(i)})-\log\cosh(kx_{r}^{(i)}m_{r}(\bm{x}^{(i)}))-\log 2\right\}$
In this paper, we study the performances of both the tensor-structure learners RISE $\hat{\bm{J}}_{r,I}$ and RPLE $\hat{\bm{J}}_{r,P}$. In Section 2, we prove rates of consistency of the RISE and the RPLE, and show explicit dependence of these rates and the minimum sample size requirement on the maximum coupling strength $\beta$ and the tensor interaction factor $k$.
1.2. Related Work on Parametric Inference in Ising Models
A closely related problem is the task of infering the inverse temperature and the external magnetic field parameters from the Ising model:
${\mathbb{P}}(\bm{x})\propto e^{\beta\bm{x}^{\top}\bm{J}\bm{x}+h\sum_{i=1}^{n}x_{i}}\quad(\bm{x}\in\{-1,1\}^{p})$
given access to a single sample from such a model. There has been a flurry of works in the past two decades in the area of parametric inference in Ising models, starting with the seminal paper due to [15], who, inspired by [7, 6], first applied the pseudolikelihood approach for parameter estimation in general spin-glass models, and showed $\sqrt{p}$-consistency of the maximum pseudolikelihood estimator in this model with $h=0$, at low temperatures. This was followed by improved results on the rate of consistency and joint estimation of $(\beta,h)$ in [8] and [21]. Recently, some of these results were also extended to tensor Ising models in [32] and [33]
1.3. Structure of the Paper
The rest of the paper is organized as follows. In Section 2, we state the main results of this paper. We begin by discussing about a general theory of $\ell^{1}$-penalized $M$-estimators from [1], that will help us prove the main results in this paper. In the same section, we use this general theory to prove theoretical guarantees for the RPLE and the RISE. In Section 3, we provide some simulation studies to demonstrate the comparative performance of these two approaches.
2. Main Results
In this section, we state and prove the main results of this paper. To begin with, we sketch a general theory about $\ell^{1}$-penalized $M$-estimators from [1], that is crucial in proving our main results,
2.1. A general theory of $\ell^{1}$-penalized $M$-estimators
We use a framework from [1] for general $\ell^{1}$-regularized $M$-estimators to establish that our method works consistently. It turns out that imposing just the following two conditions on the loss function is enough to control the error of the $\ell_{1}$-regularized $M$-estimator:
$\hat{\bm{J}}_{r}:=\arg\min_{\bm{J}_{r}\in\mathbb{R}^{{\binom{p-1}{k-1}}}}\mathcal{L}(\bm{J}_{r};\mathfrak{X}^{n})+\lambda\|\bm{J}_{r}\|_{1}$
for a general convex and differentiable loss function $\mathcal{L}$.
Condition 1.
Let $\lambda$ be the $\ell_{1}$-penalty parameter. Then, the gradient of the objective function at the true neighborhood $\bm{J}_{r}$ satisfies:
$2\|\nabla\mathcal{L}(\bm{J}_{r};\mathfrak{X}^{n})\|_{\infty}\leqslant\lambda$
Condition 1 ensures that if the maximum degree of the hypergraph with (weighted) adjacency $\bm{J}$ is $d$, then the difference $\eta_{r}:=\hat{\bm{J}_{r}}-\bm{J}_{r}$ lies within the two-sided cone:
$K:=\left\{\eta\in\mathbb{R}^{p-1\choose k-1}~{}\Big{|}~{}\|\eta\|_{1}\leqslant 4\|\eta_{S}\|_{1}\right\},$ (4)
where $S$ denotes the indices of all non-zero entries of $\bm{J}_{r}$. Note that (4) follows from Lemma 1 and Example 1 in [1].
Condition 2 (Restricted Strong Convexity).
There exists $R>0$ such that for all $\eta_{r}\in K$ with $\|\eta_{r}\|_{2}\leqslant R$, there exists a constant $\kappa>0$, such that
$\mathcal{L}(\bm{J}_{r}+\eta_{r};\mathfrak{X}^{n})-\mathcal{L}(\bm{J}_{r};\mathfrak{X}^{n})-\langle\nabla\mathcal{L}(\bm{J}_{r};\mathfrak{X}^{n}),\eta_{r}\rangle\geqslant\kappa\|\eta_{r}\|_{2}^{2}.$
The second condition guarantees that the loss function is strongly convex in a conically restricted neighborhood of ($\mathbb{R}^{p-1\choose k-1}$. We will verify Conditions 1 and 2 for the RISE and the RPLE later. With these two conditions in hand, we can finally proceed to bound the error of estimation of the neighborhood parameter $\bm{J}_{r}$. The following proposition (Theorem 1 in [1]) serves as the fundamental result behind our main theorem about the theoretical guarantee of the RPLE:
Proposition 1.
Under Conditions 1 and 2 with $R>3\sqrt{d}\lambda/\kappa$, we have:
$\|\hat{\bm{J}}_{r}-\bm{J}_{r}\|_{2}\leqslant\frac{3\lambda\sqrt{d}}{\kappa}.$
The following restricted eigenvalue condition on the tensor covariance structure is also necessary for our analysis:
Assumption 1.
Let $Q:=\mathbb{E}[\bm{X}_{\cdot r}\bm{X}_{\cdot r}^{\top}]$ where $\bm{X}_{\cdot r}:=(X_{r_{1}}\ldots X_{r_{k-1}})_{(r_{1},\ldots,r_{k-1})\in T_{r}}$. Assume that there exists a constant $\alpha>0$, such that:
$\inf_{\bm{v}\in K\setminus\{\bm{0}\}}\frac{\bm{v}^{\top}Q\bm{v}}{\|\bm{v}\|_{2}^{2}}\geq\alpha~{}.$
Note that, a sufficient condition for Assumption 1 to hold, is that the minimum eigenvalue of $Q$ is bounded below by $\alpha$. With this structure in hand, we are finally ready to prove the theoretical guarantees of the RISE and RPLE.
2.2. The Regularized Pseudo-Likelihood Estimator
In this section, we analyze the performance of the RPLE. The main result about the rate of convergence of the RPLE is stated below:
Theorem 1.
Suppose that $d$ denotes the maximum degree of the hypernetwork with adjacency $\bm{J}$, and the regularization parameter $\lambda$ is chosen as:
$\lambda:=4\sqrt{2}k!\sqrt{\frac{\log\left(4{\binom{p-1}{k-1}}/\varepsilon\right)}{n}}.$
Then, there exist constants $M_{1},M_{2}>0$, such that for every node $r\in V$ and any $\varepsilon\in(0,1)$, if
$n>M_{1}\frac{d^{2}}{\alpha^{2}}\max\left\{e^{4k!\beta d},1\right\}\log\frac{{\binom{p-1}{k-1}}}{\varepsilon},$
then the following holds with probability at least $1-\varepsilon$:
$\|\hat{\bm{J}}_{r,P}-\bm{J}_{r}\|\leqslant\frac{M_{2}\sqrt{d}e^{2k!\beta d}}{\alpha k!}\sqrt{\frac{\log\frac{{\binom{p-1}{k-1}}}{\varepsilon}}{n}}.$
Proof.
Let $\gamma\in(0,1)$ which will be chosen suitably later. In view of Lemma 1, we know that with probability at least $1-\gamma\varepsilon$, Condition 1 is satisfied with
$\lambda:=2k!\sqrt{\frac{8\log\left(2{\binom{p-1}{k-1}}/\gamma\varepsilon\right)}{n}}~{}.$
Next, note that for some $\eta_{r}\in K$, one has by the Cauchy-Schwarz inequality,
$\|\eta_{r}\|_{1}\leq 4\sqrt{d}\|\eta_{r}\|_{2}~{}.$
Define $R:=c/\sqrt{d}k!$, where $c$ is a constant to be chosen later. Then, in view of Lemma 4, Condition 2 is satisfied with
$\kappa:=\frac{\alpha(k!)^{2}e^{-2k!\beta d}}{4(1+4c)}$
with probability at least $1-(1-\gamma)\varepsilon$ as long as $n>2^{11}\frac{d^{2}}{\alpha^{2}}\log\frac{2{\binom{p-1}{k-1}}^{2}}{(1-\gamma)\varepsilon}$. Now, we can apply Proposition 1 as long as $\frac{c}{\sqrt{d}k!}>3\frac{\sqrt{d}\lambda}{\kappa}$, which is equivalent to the condition:
$n>\frac{C_{1}d^{2}}{\alpha^{2}}e^{4k!\beta d}\log\left(\frac{2{\binom{p-1}{k-1}}}{\gamma\varepsilon}\right)$
for some constant $C_{1}>0$. In that case, we have:
$\|\hat{\bm{J}}_{r,P}-\bm{J}_{r}\|\leqslant\frac{L\sqrt{d}e^{2k!\beta d}}{\alpha k!}\sqrt{\frac{\log\frac{2{\binom{p-1}{k-1}}}{\gamma\varepsilon}}{n}}$
for some constant $L>0$. Theorem 1 now follows, by taking for example, $\gamma=1/2$ and $c=1$. ∎
2.3. The Regularized Interaction Screening Estimator
In this section, we analyze the performance of the RISE. The main result about the rate of convergence of the RISE is stated below:
Theorem 2.
Suppose that $d$ denotes the maximum degree of the hypernetwork with adjacency $\bm{J}$, and the regularization parameter $\lambda$ is chosen as:
$\lambda=2\sqrt{2}k!e^{k!\beta d}\sqrt{\frac{\log\frac{4{\binom{p-1}{k-1}}}{\varepsilon}}{n}},$
Then, there exist constants $M_{1},M_{2}>0$, such that for every node $r\in V$ and any $\varepsilon\in(0,1)$, if
$n>M_{1}\frac{d^{2}}{\alpha^{2}}\max\{e^{4k!\beta d},1\}\log\frac{{\binom{p-1}{k-1}}}{\varepsilon},$
the following properties hold with probability at least $1-\varepsilon$:
$\|\hat{\bm{J}}_{r,I}-\bm{J}_{r}\|\leqslant\frac{M_{2}\sqrt{d}e^{2k!\beta d}}{\alpha k!}\sqrt{\frac{\log\frac{{\binom{p-1}{k-1}}}{\varepsilon}}{n}}.$
Proof.
In view of Lemma 5, we know that with probability at least $1-\frac{\varepsilon}{2}$, Condition 1 is satisfied with
$\lambda:=2\sqrt{2}k!e^{k!\beta d}\sqrt{\frac{\log\frac{4{\binom{p-1}{k-1}}}{\varepsilon}}{n}}$
If we set $R=\frac{2}{\sqrt{d}k!}$, then in view of Lemma 7, Condition 2 is satisfied with
$\kappa:=\frac{\alpha(k!)^{2}e^{-k!\beta d}}{20}$
with probability at least $1-\frac{\varepsilon}{2}$ as long as $n>2^{11}\frac{d^{2}}{\alpha^{2}}\log\frac{4{\binom{p-1}{k-1}}^{2}}{\varepsilon}$. Now, we can apply Proposition 1 as long as $\frac{2}{\sqrt{d}k!}>3\frac{\sqrt{d}\lambda}{\kappa}$, which is equivalent to the condition:
$n>\frac{Cd^{2}}{\alpha^{2}}e^{4k!\beta d}\log\frac{4{\binom{p-1}{k-1}}}{\varepsilon}$
for some constant $C>0$. In that case, we have:
$\|\hat{\bm{J}}_{r,I}-\bm{J}_{r}\|\leqslant\frac{D\sqrt{d}e^{2k!\beta d}}{\alpha k!}\sqrt{\frac{\log\frac{4{\binom{p-1}{k-1}}}{\varepsilon}}{n}}$
for some constant $L>0$, which completes the proof of Theorem 2. ∎
Figure 1. Plot of the estimation error (vertical axis) against the sample size $n$ (horizontal axis). The top-left window corresponds to the case $\beta=1$, the top-right window to the case $\beta=1.5$, the bottom-left window to the case $\beta=2$ and the bottom-right window to the case $\beta=2.5$.
3. Simulation Studies
In this section, we present some numerical experiments that illustrate the comparative performance of the two tensor recovery algorithms. The Julia package GraphicalModelLearning is modified and used to learn the Ising models with the interaction screening and pseudo-likelihood approaches. The performances of the RPLE and RISE are dependent on the regularization coefficient $\lambda$, which was set as $\lambda:=c\sqrt{\log(4{\binom{p-1}{k-1}}/\varepsilon)/n}$, where $c$ is a constant tuned according to the Bayesian Information Criterion (BIC). To be precise, for each node $r$, the optimal value of $\lambda$ is tuned by minimizing the BIC value:
$\mathrm{BIC}_{r}(\lambda):=\mathcal{L}_{p}(\bm{\hat{}}{J}_{r,\lambda};\mathfrak{X}^{n})+\mathrm{df}(\lambda)\log p$
over a grid of values of $\lambda$, where $\mathcal{L}_{p}$ denotes the objective function, which is $\mathcal{S}$ for the RISE and $\ell$ for the RPLE, and $\hat{J}_{r,\lambda}$ is the estimate corresponding to the penalty parameter $\lambda$.
Figure 2. Plot of the estimation error (vertical axis) against the maximum coupling intensity $\beta$ (horizontal axis) for sample size $n=10^{5}$.
For our numerical experiments, we work with the Ising model on $3-$regular, $3-$uniform hypergraphs on $16$ nodes. In Figure 1, we plot the estimation error of the recovered hypergraphs against the sample size $n$, for four values of the maximum coupling intensity $\beta$, namely $1,1.5,2,2.5$. We observe that the estimation errors are higher for both the RPLE and the RISE for higher values of $\beta$, as expected. Further, the simulations suggest that the RISE performs better than the RPLE in terms of the estimation error, which leaves open the natural question of whether this better performance of the RISE can be demonstrated theoretically. In Figure 2, we plot the estimation error for a fixed sample size $n=10^{5}$ against varying maximum coupling strengths $\beta$. The figure clearly shows an exponentially increasing dependence of the estimation error on $\beta$, as suggested in Theorems 1 and 2.
Appendix A Verification of Conditions 1 and 2 for the Pseudolikelihood
In this section, we verify conditions 1 and 2 for the pseudolikelihood, which is a crucial step in the proof of Theorem 1. We begin by verifying Condition 1.
Lemma 1.
For any $\varepsilon>0$, with probability at least $1-\varepsilon$, we have:
$\left\|\nabla\ell(\bm{J}_{r};\mathfrak{X}^{n})\right\|_{\infty}\leqslant k!\sqrt{\frac{8\log\left(2{\binom{p-1}{k-1}}/\varepsilon\right)}{n}},$
Proof.
To begin with, note that:
$\frac{\partial\ell(\bm{J}_{r};\mathfrak{X}^{n})}{\partial J_{r,r_{1},\ldots,r_{k-1}}}=-\frac{1}{n}\sum_{i=1}^{n}k!X_{r_{1}}^{(i)}\ldots X_{r_{k-1}}^{(i)}\left(X_{r}^{(i)}-\tanh\left(km_{r}(\bm{X}^{(i)})\right)\right)=:-\frac{1}{n}\sum_{i=1}^{n}Y_{i}$
where $Y_{1},\ldots,Y_{n}$ are mean 00, independent random variables bounded between $-2k!$ and $2k!$. Hence, by Hoeffding’s inequality, we have:
${\mathbb{P}}\left(\left|\frac{\partial\ell(\bm{J}_{r};\mathfrak{X}^{n})}{\partial J_{r,r_{1},\ldots,r_{k-1}}}\right|\geq\frac{{2\sqrt{2}}k!t}{\sqrt{n}}\right)\leq 2e^{-t^{2}}~{}.$
Choosing $t:=\sqrt{\log\left(2{\binom{p-1}{k-1}}/\varepsilon\right)}$, we thus have:
${\mathbb{P}}\left(\left|\frac{\partial\ell(\bm{J}_{r};\mathfrak{X}^{n})}{\partial J_{r,r_{1},\ldots,r_{k-1}}}\right|\geq k!\sqrt{\frac{8\log\left(2{\binom{p-1}{k-1}}/\varepsilon\right)}{n}}\right)\leq\frac{\varepsilon}{{\binom{p-1}{k-1}}}~{}.$
It now follows from a simple union bound, that:
${\mathbb{P}}\left(\left\|\nabla\ell(\bm{J}_{r};\mathfrak{X}^{n})\right\|_{\infty}\geq k!\sqrt{\frac{8\log\left(2{\binom{p-1}{k-1}}/\varepsilon\right)}{n}}\right)\leq\varepsilon~{}.$
The proof of Lemma 1 is now complete. ∎
Next, we verify Condition 2. Towards this, first note that for any fixed node $r\in[p]$, the Hessian matrix of the negative log pseudolikelihood is given by:
$\nabla^{2}\ell(\bm{J}_{r};\mathfrak{X}^{n})=-\frac{1}{n}\sum_{i=1}^{n}\delta_{r}(\bm{X}^{(i)};\bm{J})\bm{X}^{(i)}_{\cdot r}(\bm{X}^{(i)}_{\cdot r})^{\top},$
where
$\delta_{r}(\bm{X};\bm{J}):=\frac{4(k!)^{2}e^{2kX_{r}m_{r}(\bm{X})}}{(e^{2kX_{r}m_{r}(\bm{X})}+1)^{2}}.$
and recall that the $\binom{p-1}{k-1}$ dimensional vector $\bm{X}_{\cdot r}$ is defined as
$\bm{X}_{\cdot r}:=(X_{r_{1}}\cdots X_{r_{k-1}})_{(r_{1},\ldots,r_{k-1})\in T_{r}}~{}.$
Next, for $\eta_{r}\in\mathbb{R}^{\binom{p-1}{k-1}}$, define the remainder of the first order Taylor expansion of $\ell(\bm{J}_{r}+\eta_{r};\mathfrak{X}^{n})$ as:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n}):=\ell(\bm{J}_{r}+\eta_{r};\mathfrak{X}^{n})-\ell(\bm{J}_{r};\mathfrak{X}^{n})-\langle\nabla\ell(\bm{J}_{r};\mathfrak{X}^{n}),\eta_{r}\rangle.$
In the next lemma, we lower bound this remainder term.
Lemma 2.
For all $\eta_{r}\in\mathbb{R}^{{\binom{p-1}{k-1}}}$, the remainder of the first order Taylor expansion is bounded below by:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{e^{-2k!\beta d}(k!)^{2}}{2k!\|\eta_{r}\|_{1}+2}~{}\eta_{r}^{\top}\hat{Q}\eta_{r}$
where $\hat{Q}:=\frac{1}{n}\sum_{i=1}^{n}\bm{X}^{(i)}_{\cdot r}(\bm{X}^{(i)}_{\cdot r})^{\top}$.
Proof.
To begin with, note that:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})=\frac{1}{n}\sum_{i=1}^{n}\left[\log\left(\frac{\cosh(km_{r}(\bm{X}^{(i)})+k!\eta_{r}^{\top}\bm{X}_{\cdot r}^{(i)})}{\cosh(km_{r}(\bm{X}^{(i)}))}\right)-k!\eta_{r}^{\top}\bm{X}_{\cdot r}^{(i)}\tanh(km_{r}(\bm{X}^{(i)}))\right]$
Using Lemma 8 with $x:=km_{r}(\bm{X}^{(i)})$ and $a:=k!\eta_{r}^{\top}\bm{X}_{\cdot r}^{(i)}$, and the inequality
$1-\tanh^{2}(x)\geq e^{-2|x|}~{},$
we have:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geq\frac{\exp\left(-2\left|km_{r}(\bm{X}^{(i)})\right|\right)(k!)^{2}}{2+2k!\|\eta_{r}\|_{1}}~{}\eta_{r}^{\top}\hat{Q}\eta_{r}~{}.$
Lemma 2 now follows on observing that $|km_{r}(\bm{X})|\leqslant k!\beta d$. ∎
In the following lemma, we provide error bounds for estimating $Q$ by $\hat{Q}$.
Lemma 3.
For every $\delta>0$, $\varepsilon>0$ and $n\geq\frac{2}{\delta^{2}}\log\frac{2\binom{p-1}{k-1}^{2}}{\varepsilon}$, we have the following with probability at least $1-\varepsilon$:
$\max_{(t_{r},t_{s})\in T_{r}^{2}}|\hat{Q}_{t_{r},t_{s}}-Q_{t_{r},t_{s}}|<\delta,$
Proof.
Note that for $t_{r}:=(r_{1},\ldots,r_{k-1})$ and $t_{s}:=(s_{1},\ldots,s_{k})$, we have
$\hat{Q}_{t_{r},t_{s}}=\frac{1}{n}\sum_{i=1}^{n}x_{r_{1}}...x_{r_{k-1}}x_{s_{1}}...x_{s_{k-1}},$
Since $|x_{r_{1}}...x_{r_{k-1}}x_{s_{1}}...x_{s_{k-1}}|\leqslant 1$, by Hoeffding’s inequality we have
$P\left(|\hat{Q}_{t_{r}t_{s}}-Q_{t_{r}t_{s}}|\geqslant\delta\right)\leqslant 2\exp\left(-\frac{n\delta^{2}}{2}\right),$
Using a union bound over the upper triangle of the symmetric matrix we have
$P\left(\max_{(t_{r},t_{s})\in T_{r}^{2}}|\hat{Q}_{t_{r}t_{s}}-Q_{t_{r}t_{s}}|<\delta\right)\geqslant 1-2{\binom{p-1}{k-1}}^{2}\exp\left(-\frac{n\delta^{2}}{2}\right)\geq 1-\varepsilon$
thereby completing the proof of Lemma 3. ∎
The following lemma provides the final step towards verifying Condition 2 for the pseudolikelihood:
Lemma 4.
For all $\varepsilon>0$, $R>0$, and $n>2^{11}\frac{d^{2}}{\alpha^{2}}\log\frac{2{\binom{p-1}{k-1}}^{2}}{\varepsilon}$, with probability at least $1-\varepsilon$, the restricted strong convexity condition holds, i.e.
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{\alpha(k!)^{2}e^{-2k!\beta d}}{4(1+4k!\sqrt{d}R)}\|\eta_{r}\|_{2}^{2},$
for all $\eta_{r}\in K$ such that $\|\eta_{r}\|_{2}\leqslant R$.
Proof.
From Lemma 2, we have:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{e^{-2k!\beta d}}{2k!\|\eta_{r}\|_{1}+2}(k!)^{2}\eta_{r}^{\top}\hat{Q}\eta_{r}.$
Since $\|\eta_{r}\|_{1}\leqslant 4\sqrt{d}\|\eta_{r}\|_{2}\leqslant 4\sqrt{d}R$, we conclude that:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{e^{-2k!\beta d}}{8k!\sqrt{d}R+2}(k!)^{2}\eta_{r}^{\top}\hat{Q}\eta_{r},$
Next, we have by Assumption 1:
$\eta_{r}^{\top}\hat{Q}\eta_{r}$ $=\eta_{r}^{\top}Q\eta_{r}+\eta_{r}^{\top}(\hat{Q}-Q)\eta_{r}$
$\geqslant\alpha\|\eta_{r}\|_{2}^{2}+\eta_{r}^{\top}(\hat{Q}-Q)\eta_{r},$
Using Lemma 3 with $\delta:=\frac{\alpha}{32d}$, we have the following for $n>2^{11}\frac{d^{2}}{\alpha^{2}}\log\frac{2{\binom{p-1}{k-1}}^{2}}{\varepsilon}$ with probability at least $1-\varepsilon$,
$\eta_{r}^{\top}(\hat{Q}-Q)\eta_{r}$ $\geqslant-\frac{\alpha}{32d}\|\eta_{r}\|_{1}^{2}$
$\geqslant-\frac{\alpha}{2}\|\eta_{r}\|_{2}^{2},$
Hence, we have:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{\alpha(k!)^{2}e^{-2k!\beta d}}{4(1+4k!\sqrt{d}R)}\|\eta_{r}\|_{2}^{2}$
which completes the proof of Lemma 4. ∎
Appendix B Verification of Conditions 1 and 2 for the Interaction Screening Approach
In this section, we verify conditions 1 and 2 for the interaction screening function, which is a fundamental step for proving Theorem 2. We start by checking Condition 1.
Lemma 5.
For any $\varepsilon>0$, with probability at least $1-\varepsilon$, we have:
$\left\|\nabla\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})\right\|_{\infty}\leqslant\sqrt{2}k!e^{k!\beta d}\sqrt{\frac{\log\frac{2{\binom{p-1}{k-1}}}{\epsilon}}{n}},$
Proof.
First, note that we have:
$\frac{\partial\mathcal{S}(\bm{J}_{r};x^{(i)})}{\partial J_{r,r_{1},\ldots,r_{k-1}}}=-\frac{1}{n}\sum_{i=1}^{n}k!X_{r_{1}}^{(i)}\ldots X_{r_{k-1}}^{(i)}X_{r}^{(i)}e^{-kX_{r}^{(i)}m_{r}(\bm{X}^{(i)})}:=-\frac{1}{n}\sum_{i=1}^{n}V_{i}$
where
$|V_{i}|\leqslant k!e^{k!\beta d},$
Since $V_{1},\ldots,V_{n}$ are i.i.d. with mean 00, we have the following by Hoeffding’s inequality:
$P\left(\left|\frac{\partial\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})}{\partial\bm{J}_{r,r_{1},...,r_{k-1}}}\right|\geqslant\frac{\sqrt{2}k!e^{k!\beta d}t}{\sqrt{n}}\right)\leqslant 2e^{-t^{2}},$
As there are ${\binom{p-1}{k-1}}$ components of the gradient vector, we can choose $t:=\sqrt{\log\left(2{\binom{p-1}{k-1}}/\epsilon\right)}$. Using a union bound we get the following with probability at least $1-\epsilon$:
$\left\|\nabla\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})\right\|_{\infty}\leqslant\sqrt{2}k!e^{k!\beta d}\sqrt{\frac{\log\frac{2{\binom{p-1}{k-1}}}{\epsilon}}{n}}~{},$
thereby completing the proof of Lemma 5. ∎
Next we check condition 2. Towards this, first note that for any fixed node $r\in V$, we can express the Hessian of the interaction screening function as:
$\nabla^{2}\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})=\frac{1}{n}\sum_{i=1}^{n}\gamma_{r}(\bm{X}^{(i)};\bm{J}_{r})\bm{X}^{(i)}_{\cdot r}(\bm{X}^{(i)}_{\cdot r})^{\top},$
where
$\gamma_{r}(\bm{X};\bm{J}_{r}):=(k!)^{2}\exp(-kX_{r}m_{r}(\bm{X})),$
Using the inequality $\gamma_{r}(\bm{X};\bm{J})\geqslant(k!)^{2}e^{-\beta d(k-1)!}$, the Hessian is lower bounded in the positive semi-definite sense by:
$\nabla^{2}\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})\succcurlyeq(k!)^{2}e^{-\beta dk!}\hat{Q}$
where recall that $\hat{Q}=\frac{1}{n}\sum_{i=1}^{n}\bm{X}^{(i)}_{\cdot r}(\bm{X}^{(i)}_{\cdot r})^{\top}$.
As before, for every $\eta_{r}\in\mathbb{R}^{\binom{p-1}{k-1}}$, define the remainder term:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n}):=\mathcal{S}(\bm{J}_{r}+\eta_{r};\mathfrak{X}^{n})-\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n})-\langle\nabla\mathcal{S}(\bm{J}_{r};\mathfrak{X}^{n}),\eta_{r}\rangle.$
In the next lemma, we lower bound this remainder term.
Lemma 6.
For all $\eta_{r}\in\mathbb{R}^{{\binom{p-1}{k-1}}}$, the remainder of the first order Taylor expansion is bounded below by:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{(k!)^{2}e^{-k!\beta d}}{2+k!\|\eta_{r}\|_{1}}\eta_{r}^{\top}\hat{Q}\eta_{r}.$
Proof.
To begin with, note that:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})=\frac{1}{n}\sum_{i=1}^{n}\left[e^{-kX_{r}^{(i)}m_{r}(\bm{X}^{(i)})}\left(e^{-k!\eta_{r}^{\top}\bm{X}_{\cdot r}^{(i)}X_{r}^{(i)}}-1+k!\eta_{r}^{\top}X_{r}^{(i)}\bm{X}_{\cdot r}^{(i)}\right)\right],$
Since $|km_{r}(\bm{X})|\leqslant k!\beta d$, we have:
$e^{-kX_{r}^{(i)}m_{r}(\bm{X}^{(i)})}\geqslant e^{-k!\beta d},$
Next, using Lemma 9 together with the inequality
$\left|k!\eta_{r}^{\top}X_{r}^{(i)}\bm{X}_{\cdot r}^{(i)}\right|\leqslant k!\|\eta_{r}\|_{1},$
we have
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})$ $\geqslant\frac{e^{-k!\beta d}}{2+k!\|\eta_{r}\|_{1}}\frac{1}{n}\sum_{i=1}^{n}\left(k!X_{r}^{(i)}\eta_{r}^{\top}\bm{X}_{\cdot r}^{(i)}\right)^{2}$
$=\frac{(k!)^{2}e^{-k!\beta d}}{2+k!\|\eta_{r}\|_{1}}\eta_{r}^{\top}\hat{Q}\eta_{r},$
The following lemma provides the final step towards verifying Condition 2 for the interaction screening approach:
Lemma 7.
For all $\varepsilon>0$, $R>0$, and $n>2^{11}\frac{d^{2}}{\alpha^{2}}\log\frac{2{\binom{p-1}{k-1}}^{2}}{\varepsilon}$, with probability at least $1-\varepsilon$, the restricted strong convexity condition holds, ie
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{\alpha(k!)^{2}e^{-k!\beta d}}{4(1+2k!\sqrt{d}R)}\|\eta_{r}\|_{2}^{2},$
for all $\eta_{r}\in K$ such that $\|\eta_{r}\|_{2}\leqslant R$.
Proof.
From Lemma 6, we have
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{(k!)^{2}e^{-k!\beta d}}{2+k!\|\eta_{r}\|_{1}}\eta_{r}^{\top}\hat{Q}\eta_{r}.$
Since $\|\eta_{r}\|_{1}\leqslant 4\sqrt{d}\|\eta_{r}\|_{2}\leqslant 4\sqrt{d}R$, we have:
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{(k!)^{2}e^{-k!\beta d}}{4k!\sqrt{d}R+2}\eta_{r}^{\top}\hat{Q}\eta_{r},$
Next, we have by Assumption 1:
$\eta_{r}^{\top}\hat{Q}\eta_{r}$ $=\eta_{r}^{\top}Q\eta_{r}+\eta_{r}^{\top}(\hat{Q}-Q)\eta_{r}$
$\geqslant\alpha\|\eta_{r}\|_{2}^{2}+\eta_{r}^{\top}(\hat{Q}-Q)\eta_{r},$
Using Lemma 3, we set $\delta=\frac{\alpha}{32d}$, we get for $n>2^{11}\frac{d^{2}}{\alpha^{2}}\log\frac{2{\binom{p-1}{k-1}}^{2}}{\varepsilon}$ with probability at least $1-\varepsilon$,
$\eta_{r}^{\top}(\hat{Q}-Q)\eta_{r}$ $\geqslant-\frac{\alpha}{32d}\|\eta_{r}\|_{1}^{2}$
$\geqslant-\frac{\alpha}{2}\|\eta_{r}\|_{2}^{2},$
Thus
$\Delta(\bm{J}_{r};\mathfrak{X}^{n})\geqslant\frac{\alpha(k!)^{2}e^{-k!\beta d}}{4(1+2k!\sqrt{d}R)}\|\eta_{r}\|_{2}^{2}.$
This completes the proof of Lemma 7. ∎
Appendix C Technical Lemmas
In this section, we prove some technical lemmas necessary for the proofs of the main results.
Lemma 8.
For every $x,a\in\mathbb{R}$, we have:
$\log\left(\frac{\cosh(x+a)}{\cosh(x)}\right)-a\tanh(x)\geqslant\frac{(1-\tanh^{2}(x))a^{2}}{2+2|a|},$
Proof.
Fixing $x\in\mathbb{R}$, let
$f(a):=\log\left(\frac{\cosh(x+a)}{\cosh(x)}\right)-a\tanh(x)-\frac{\mathrm{sech}^{2}(x)a^{2}}{2+2|a|},$
First, consider the case $a>0$, and note that:
$f^{\prime}(a)=\tanh(x+a)-\tanh(x)-\frac{1}{2}\mathrm{sech}^{2}(x)\left(1-\frac{1}{(a+1)^{2}}\right),$
and
$f^{\prime\prime}(a)=\mathrm{sech}^{2}(x+a)-\mathrm{sech}^{2}(x)\cdot\frac{1}{(a+1)^{3}},$
We claim that there exists $\delta>0$ such that $f^{\prime\prime}>0$ on $(0,\delta)$ and $f^{\prime\prime}<0$ on $(\delta,\infty)$. Given this claim, we know that since $f^{\prime}(0)=0$ and $\lim_{a\rightarrow\infty}f^{\prime}(a)=\frac{1}{2}(\tanh(x)-1)^{2}>0$, we must have $f^{\prime}(a)>0$ for all $a>0$. Since $f(0)=0$, we must thus have $f(a)>0$ for all $a>0$, which proves Lemma 8 for all $x\in\mathbb{R}$ and $a\geq 0$. If $a<0$, then note that we can apply this result with $x$ replaced by $-x$ and $a$ replaced by $-a$, to complete the proof of Lemma 8. Hence, it only suffices to prove the claim, in order to complete the proof of Lemma 8. Towards this, define $g(a):=\cosh^{2}(x)\cdot(a+1)^{3}-\cosh^{2}(x+a)$, whence
$g^{\prime}(a)$ $=3\cosh^{2}(x)(a+1)^{2}-\sinh(2x+2a)$
$g^{\prime\prime}(a)$ $=6\cosh^{2}(x)(a+1)-2\cosh(2x+2a)$
$g^{(3)}(a)$ $=6\cosh^{2}(x)-4\sinh(2x+2a),$
Note that $g^{(3)}$ is strictly decreasing and $\lim_{a\rightarrow\infty}g^{(3)}(a)=-\infty$. Meanwhile, $g^{\prime\prime}(0)=2\cosh^{2}(x)+2>0$ and $\lim_{a\rightarrow\infty}g^{\prime\prime}(a)=-\infty$. Hence, $g^{\prime\prime}$ is positive on $(0,K)$ and negative on $(K,\infty)$ for some $K>0$. Next, since $g^{\prime}(0)=\cosh(x)(2e^{-x}+\cosh(x))>0$ and $\lim_{a\rightarrow\infty}g^{\prime}(a)=-\infty$, the previous argument implies that $g^{\prime}$ is positive on $(0,L)$ and negative on $(L,\infty)$ for some $L>0$. Finally, since $g(0)=0$ and $\lim_{a\rightarrow\infty}g(a)=-\infty$, there exists $\delta>0$, such that for $0<a<\delta$, $\cosh^{2}(x)\cdot(a+1)^{3}>\cosh^{2}(x+a)$, and for $a>\delta$, $\cosh^{2}(x)\cdot(a+1)^{3}<\cosh^{2}(x+a)$. So for $0<a<\delta$, $f^{\prime\prime}(a)>0$, and for $a>\delta$, $f^{\prime\prime}(a)<0$. This proves the claim, and completes the proof of Lemma 8. ∎
Lemma 9.
For all $x\in\mathbb{R}$, we have:
$e^{-x}-1+x\geq\frac{x^{2}}{2+|x|}\quad\text{for all}~{}x\in\mathbb{R}$
Proof.
Define:
$g(x)=e^{-x}-1+x-\frac{x^{2}}{2+|x|}~{}.$
Then, for $x>0$, we have:
$g^{\prime}(x)=-e^{-x}+\frac{4}{(2+x)^{2}}=\left(1+\frac{x}{2}\right)^{-2}-(e^{\frac{x}{2}})^{-2}>0~{}.$
Since $g(0)=0$, we must thus have $g(x)>0$ for all $x>0$, thereby proving Lemma 9 for $x>0$. Now, suppose that $x<0$. Since $\sinh(-x)>-x$, we have:
$e^{-x}-1+x>e^{x}-1-x\geq\frac{x^{2}}{2+|x|}$
where the last inequality follows on applying Lemma 9 on $-x>0$. ∎